OpenStudy (anonymous):

Is this right? 20. Determine the solution of the system. Show your work. Y = 5x – 9 -5x -5 -14 Y = 2x + 6 -2 -2 4

OpenStudy (anonymous):

no. these equations are already solved for y, so set them equal to each other to solve for x. 5x - 9 = 2x + 6

OpenStudy (anonymous):

Oh, thanks. thought it would of been harder then that XD

OpenStudy (anonymous):

Can you help me out on my next problem?

OpenStudy (anonymous):

ok. did you get a y-value for the one above?

OpenStudy (anonymous):

Isn't that the same as the x vaule?

OpenStudy (anonymous):

5x -9 =2x +6

OpenStudy (anonymous):

not necessarily. when they say "solve the system" they're asking for a point that makes both equations true. that point will have an x-coordinate and a y-coordinate. Once you have the x, substitute it into either one to get y

OpenStudy (anonymous):

what did you get when you solved that equation for x?

OpenStudy (anonymous):

-4x

OpenStudy (anonymous):

5x -9 =2x +6 I added both x terms then subtracted -9 from 6 7 x- 3 7/7 -3/7

OpenStudy (anonymous):

You can't add/subtract things on different sides of an equal sign. You can move things from one side to the other using addition or subtraction. Let's start by getting all the terms with x on the left side. To do that we need to eliminate 2x from the right side. Since 2x is positive, subtract 2x from both sides.